OpenStudy (anonymous):

Find all zeroes for f(x)=x^4-3x^2-4

OpenStudy (anonymous):

Zeros, sorry!

OpenStudy (mathstudent55):

Use a substitution, u = x^2, then u^4 = x^2 x^4 - 3x^2 - 4 = 0 Apply substitution: u^2 - 3u - 4 = 0 Now factor it the normal way: (u -4)(u + 1) = 0 u - 4 = 0 or u + 1 = 0 u = 4 or u = -1 But since u = x^2, x^2 = 4 or x^2 = -1 x = +-2 or x = +-i

OpenStudy (anonymous):

@mathstudent55 thank you so much!!! That really helped a lot :) I was taught a more confusing way, but I love this new method. Thanks!

OpenStudy (anonymous):

@mathstudent55 thank you so much!!! That really helped a lot :) I was taught a more confusing way, but I love this new method. Thanks!

OpenStudy (mathstudent55):

You're very welcome!